Saddle Inc. has two types of handbags: standard and custom. The controller has decided to use a plantwide overhead rate based on direct labor costs. The president has heard of activity-based costing and wants to see how the results would differ if this system were used. Two activity cost pools were developed: machining and machine setup. Presented below is information related to the company’s operations.

Standard

Custom

Direct labor costs $50,000 $100,000
Machine hours 1,200 1,500
Setup hours 90 390


Total estimated overhead costs are $285,000. Overhead cost allocated to the machining activity cost pool is $189,000, and $96,000 is allocated to the machine setup activity cost pool

Determine the difference in allocation between the two approaches.

Traditional costing
Standard

$enter a dollar amount

Custom

$enter a dollar amount

Activity-based costing
Standard

$enter a dollar amount

Custom

$enter a dollar amount
Answer #1

Solution:

Traditional Costing:

Predetermined Overhead rate = Estimated Overhead / Total direct labor costs = $285000 / ($50000 + $100000) = $285000 / $150000 = 190%

Overhead allocated to Standard = $50000 *190% = $95,000

Overhead allocated to Custom = $100000 *190% = $190,000

Activity based costing:

Activity Cost Pool Overhead Cost Expected Activity Activity Rate
Machining $1,89,000 2700 Machine Hours $70.00 per Mchine hour
Machine Setup $96,000 480 Setup hours $200.00 per setup hour
Overhead allocated to Standard
Activity Cost Pool Activity Rate Activity Usage Overhead Allocated
Machining $70.00 1200 84000
Machine Setup $200.00 90 18000
Overhead allocated to Standard 102000
Overhead allocated to Custom
Activity Cost Pool Activity Rate Activity Usage Overhead Allocated
Machining $70.00 1500 105000
Machine Setup $200.00 390 78000
Overhead allocated to Custom 183000